What is the equilibrium Zn2+ concentration when 1.36 L of a 0.166 M zinc sulfate solution are mixed with 1.43 L of a 0.286 M ammonium sulfide solution?

[Zn2+] =  M
Answer #1

The reaction of zinc sulfate (ZnSO4) and ammonium sulfide (NH4)2S is

ZnSO4(aq) + (NH4)2S(aq) ----------> ZnS(s) + (NH4)2SO4(aq)

This equation indicates that  when we react 1 mole aqueous ZnSO4 adn 1 mole (NH4)2S, it produces 1 mole insoluble ZnS and 1 mole aqueous (NH4)2SO4

Now the ZnSO4 present in the reactant side , being soluble in water can be converted to Zn+2. But the ZnS present in the product, being insoluble cant be converted into Zn+2

Thus after the complete reaction, the ZnSO4 left will only produce Zn+2

Now given ZnSO4 taken = 1.36 L of a 0.166 M

And (NH4)2S taken = 1.43 L of a 0.286 M

Now after mixing the two, the final volume = 1.36 L + 1.43 L = 2.79‬ L

Thus the new initial concentration of ZnSO4 = M1V1/V2 = 0.166 M * 1.36 L / 2.79‬ L = 0.321 M

Similarly new initial concentration of (NH4)2S = M1V1/V2 = 0.286 M * 1.43 L / 2.79‬ L = 0.146 M

Thus the concentration at equilibrium is

Reaction ZnSO4 (NH4)2S ZnS(s) + (NH4)2SO4(aq)
Initial 0.321 M 0.146 M 0 + 0
Change -  0.146 M - 0.146 M +0.146 M  + 0.146 M
Equilibrium 0.175 0 0.146 M  + 0.146 M

That is after the complition of reaction, amount of ZnSO4 left at equilibrium = 0.175 M

Now this ZnSO4 will dissociate as   ZnSO4 -----------> Zn+2 + SO42-

It indicates that from 1 mole of ZnSO4, we get 1 mole of Zn+2

Thus form 0.175 M of ZnSO4, we get 0.175 M of Zn+2

Hence finally at equilibrium, [Zn2+] =  0.175 M